# Experiential Explanations for Reinforcement Learning

### v4

@@ 4 Evaluation overview
Our evaluations focus on exploring how beneficial experiential evaluations are to people, and how people use explanations to reason about their own decisions through quantitative and qualitative analyses. Our human evaluations address the following research questions:
@@
    Do Experiential Explanations improve the users’ predictions of the agent’s actions over baselines?
@@
    Do users find experiential explanations satisfactory compared to baselines?
@@
    What features of the explanations do users perceive as helpful in predicting agent actions?
@@
    Are the explanations effective in more complex domains?
We also examined the faithfulness (RQ 5) of our explanation system automatically by looking at how closely the influence predictor’s values estimate the agent’s actions.
In the following sections, we detail the two environments on which we evaluated our explanations. The first is the MiniGrid environment, where we examine the performance of a global version of the explanations in a sparse-reward setting with two reward classes. In the second study, we use the Crafter environment to explore the effectiveness of the local explanations generated by our method in a more complex environment with sequential multi-step goals, along with more granular reward classes. Moreover, we want to see if the local Experiential Explanations can convey these complex values to the participants coherently.
@@ 5 Experiment 1: MiniGrid Environment
In this study, we evaluate our explanations in an environment with two simple reward sources and simplified grid-navigation settings, with only one possible counterfactual action at the decision point (see Figure 3). Our aim is to answer our research questions (RQ1, RQ2, and RQ3) where we evaluate the performance of our explanation method with quantitative and qualitative analyses. We additionally automatically validate the faithfulness of the explanations through automated tests (RQ5). We built our experimental domain on MiniGrid [28], a popular minimalistic grid world environment that offers a variety of tasks to test RL algorithms.
@@ 5.1 Model Setup and Environment Details
    Negative rewards heavily influence the counterfactual trajectory. “If the agent goes down, it will pass through regions influenced by the dangerous blue obstacles; going up feels safer”.
@@
    Goal rewards have a lower influence on the counterfactual trajectory. “If the agent goes right, it will pass through regions less influenced by the green goal; going left is better”.
@@ 5.2 Baselines
In this experiment, we focus on similar global baseline explanations that the agent can generate. We compared our method with the following baselines:
@@
5.3 Method
@@
We designed our human-participant evaluation following the framework proposed by Hoffman et al. [31], evaluating our system for understandability, performance, and user satisfaction. We conducted our online evaluation between subjects using Prolific [^3] http://prolific.co with 81 participants aged 21-71 years old (M=37.41, SD=11.03); 39.5% of the participants identified as women. The average study completion time was approximately 10 minutes, and we compensated each participant with \$3.75. The study consists of two parts: answering a series of prediction tasks (RQ1, RQ3) and a satisfaction survey (RQ2, RQ3).
In the first part, we presented three episodes to each participant, as shown in Figure 3. During each episode, the agent—the red triangle—must navigate to the green square. Some colored shapes are associated with negative rewards, but the participant is not told which ones. After each episode, the participant is asked to predict whether the agent will go up or down. The participant has no information at the beginning of the first episode to make an informed choice, and the participant choice is random—we do not evaluate on this choice. Further, upon the participant’s choice, we assign which object is dangerous such that the participant is always wrong. This established the conditions for which the participant must now learn to correct their understanding of the agent’s interactions with the environment. We use that assignment of dangerous object consistently throughout the second and third episodes. The participant tries to predict the agent’s choice after the second and third episodes (without the object assignment switching); we expect to see the participants in the experimental condition to have higher accuracy rates than in the baseline conditions.
After each episode, participants fill out a brief text entry question explaining the rationale behind their choice.
The second part of the study was a satisfaction survey consisting of five-point Likert scale questions asking the user to rate understandability, satisfaction, amount of details, completeness, usefulness, precision, and trust for the explanations they saw. It also had an open-ended response question asking how the explanation helped or did not help them understand the agent. Appendix D includes more details about the explanations and survey questions.
@@ 5.4 Quantitative Results
RQ1: Prediction Correctness. We measure whether the participant can correctly predict the agent trajectory after interacting with the second and third episodes (Figure 3).
@@
RQ2: Satisfaction Survey. For each type of explanation, we analyzed the ratings of the participants for the Likert scale questions on the satisfaction survey. Experiential Explanations had the highest average scores on all dimensions, except trust. A one-way analysis of variance (ANOVA) followed by Tukey’s honestly significant difference test revealed that the Experiential Explanations $(M=3.62,SD=1.32)$ were preferred over the Heatmaps $(M=2.42,SD=1.35)$ in completeness ($F(3,81)=3.032,p<0.05$). Experiential Explanations $(M=3.71,SD=1.0)$ were preferred over the Heatmaps $(M=2.63,SD=1.38)$ in satisfaction($F(3,81)=3.106,p<0.05$). Experiential Explanations $(M=4.38,SD=0.59)$ were preferred over Q-Value Explanations $(M=3.28,SD=1.01)$ and Heatmap Explanations$(M=3.0,SD=1.37)$ in the dimension of understandability ($F(3,81)=5.572,p<0.05$).
5.5 Qualitative Analysis
@@
To address RQ3, we thematically analyzed open response questions for the tasks and satisfaction survey to better understand how the participants used the explanations. Thematic analysis is one of the most common qualitative research methods and involves identifying, analyzing, and reporting patterned responses or meanings (i.e. themes[32]. It is most appropriate for understanding a set of experiences, thoughts, or behaviors across data [33]. To conduct the thematic analysis for each of these codes, three authors performed the following established process, as outlined by [32]: (1) We familiarized ourselves with the data through repeated active readings. (2) We developed a coding framework consisting of a set of codes organizing the data at a granular level and the corresponding well-defined and distinct definitions. (3) Each of the authors independently coded the entirety of the data. If a code was unclear or a new code was needed, the authors consulted with each other to modify the coding framework. If a change was made, each author independently re-analyzed the entirety of the data. (4) We jointly examined the codes to derive themes and findings. We compared our application of the codes for each response to ensure consistency. Discrepancies were counted and used to calculate the inter-rater reliability. The process iterated until all discrepancies were resolved. We identified three themes: Usefulness, Format, and Utilization.
@@
Explanation Usefulness. The theme of Usefulness emerged in the open-ended comments about how participants perceived the explanations. Because these comments were made alongside users’ explanations of their reasoning, this provides additional information on how useful users found the explanations. We identified three Usefulness codes: Useful, Not Useful, and Unclear. We coded a response Useful if a participant explicitly mentioned that the explanation was useful, for example, “I understood that purple was dangerous and that the agent should avoid getting near it (EE8).” Responses were coded Not Useful if a participant explicitly stated that the explanation was not useful, for example, “I tried to read the chart but found it confusing and unclear. I didn’t understand it very well (HE10).”. Responses were coded Unclear if the user’s perception was ambiguous, for instance, “I found the explanations a bit too brief. For instance, what makes the purple circles feel safer than the blue squares (EE7)?” The initial inter-rater reliability for Usefulness was 81.5%.
@@
Explanation Format. Participant responses were also coded for Format. We derived three codes for Format: Need More Information, Hard to Understand, and Suggestion. Responses coded as Need More Information explicitly stated why the explanation was lacking and what data was desired. If the response said that the explanation was not clear, it was coded as Hard to Understand, and if the response offered another way to explain, it was coded as Suggestion. Only responses from individuals who received explanations and referenced the explanation’s content could be coded for Format (41.98% of our dataset). Participants often expressed a Need for More Information when shown Heatmap explanations, followed by Q-value explanations. Participants who did not find the explanations useful had more Format comments. The initial inter-rater reliability was 76.47%.
@@
We compared the frequency with which utilization codes appeared in open-responses. If a participant used the same types of information as the explanation provided, we considered the participant as “Utilizing” the explanation. Experiential Explanations provide Object and Reward Related information. Heatmap and Q-Value Explanations provide Reward-Related Explanations. No Explanation only provides for Prior Behavior. 71% of the participants who were shown Experiential Explanations utilized the exact kinds of information presented in the explanation, compared to 21% and 28% for the Heatmap and Q-Value explanations, respectively. Although the Q-Value and Heatmap explanations both presented Reward-Related information, participants still relied heavily on Prior Behavior to understand the agent. Experiential Explanations were 13 and 15 times more likely to have the information they provided be Utilized and Useful with respect to Heatmap and Q-Value explanations, respectively. We also found that among those who used components of explanations in their reasoning, 55.88% answered correctly on both questions ($p<0.05,df=1,\chi^{2}=3.849$). But within the Experiential Explanations group, the percentage was 84.62% and that group was more statistically significant to get both questions correctly than other baseline explanation groups ($p<0.05,df=3,\chi^{2}=12.22$).
@@
Figure 4 shows the distribution of the reasons the participants referenced for each type of explanation. Participants who saw Experiential Explanations most used Object-related and Reward-Related information. Participants who saw Q-Value explanations referenced Goal-Related reasons the most. Participants who saw no explanation most often utilized Prior Behavior, followed by Object-Related reasons. Notably, participants who saw Experiential Explanations were 4 times less likely to rely on Prior Behavior, indicating they did not rely as much on watching the agent to understand it. Those who saw Heatmap and Q-Value Explanations relied heavily on Prior Behavior and Utilization patterns mimicking those of the No Explanation condition.
5.6 Faithfulness Evaluation
To address RQ5, a key consideration of explanation generation is the extent to which the explanations are faithful to the underlying system [34]. Influence predictors sit “outside of the black-box”, which is the policy model that drives the agent. However, influence predictors are trained alongside the main policy model, and thus, if they learned correctly, should be able to estimate the agent’s actions with a high degree of accuracy. We don’t expect them to replicate the agent’s actions as they are learning influences instead of generating actions. To assess the faithfulness of influence predictors, we evaluated the accuracy of influence predictors’ combined decision in predicting the agents.
To calculate the faithfulness, we combine the predictions of the influence models using two methods. The first is by taking the $action=\operatorname*{argmax}_{a}(\sum_{i}P_{i}(s,a)-\sum_{j}N_{j}(s,a))$ where $P_{i}$ is a positive influence predictor and $N_{j}$ is a negative influence predictor. The second method is to train a classifier on the predicted positive and negative influence values as features and the agent’s actions as the ground truth. We used an SVM classifier with an RBF kernel. We also examined how the predictors’ accuracy changed in critical regions; when influences were stronger than specified influence thresholds.
For our first test, the accuracy of the influence models in predicting the agent’s actions, we found the accuracy of the prediction of the direct aggregation is $72\%$ and the accuracy of the classifier $76\%$. In certain critical states where at least one influence predictor exceeded a threshold of $0.100$, the aggregate of influence predictors agreed with the main policy $83\%$, covering at least $20\%$ of the environment. At higher thresholds, agreement jumps above $90\%$ indicating that when decisions are the most important, faithfulness is very high. Looking at the positive influence predictor alone, at a threshold of $0.200$ and higher, it alone can predict the agent with a $100\%$ accuracy. This is to be expected as the influence predictors cannot dictate what states and actions get explored—it is controlled by the main policy and the main training algorithm—so states that are less critical to positive reward attainment are under-sampled and prone to more error in both the main policy and the influence predictors.

### v5

@@ 4 Evaluation overview
Our evaluations focus on exploring how beneficial Experiential Explanations are to participants in understanding RL agents and how they use these explanations to reason about their own decisions through quantitative and qualitative analyzes. We conducted two studies to evaluate our explanations with human participants in simple and complex environments. Our simple environment is a grid world built on Minigrid [29] environment. Our complex environment is the Crafter game [30], a 2D version of the popular commercial game, Minecraft.
Our evaluations address the following research questions in both settings.
@@
    Do explanations derived from influence predictors accurately represent the actual values of the agent?
@@
    How do Experiential Explanations improve users’ ability to predict the agent’s actions compared to baselines?
@@
    How does user satisfaction with Experiential Explanations compare to baseline explanations?
@@
    How do users who got Experiential Explanations compare in their reasoning with those who received baseline explanations?
In both simple and complex environments, we first address RQ1, which focuses on examining the faithfulness of the explanation, a key consideration of explanation generation, which measures the extent to which the explanations are faithful to the underlying system [31]. Influence predictors sit “outside of the black-box”, whereas the main policy model that drives the agent is “inside”. However, influence predictors are trained alongside the main policy model, and thus, if they learned correctly, should be able to estimate the agent’s actions with a high degree of accuracy. We don’t expect them to replicate the agent’s actions as they are learning influences instead of generating actions. To assess the faithfulness of influence predictors, we evaluated the accuracy of the combined decision of influence predictors in predicting the agents’ actions and how closely the combined influence predictor’s values estimate the agent’s.
We then designed and conducted a our human-participant study following the framework proposed by Hoffman et al. [32], evaluating our explanations for actionability, satisfaction, and utilization. We examine the actionability of the given explanations through a prediction and ordering tasks of the agent’s actions and sub-goals (RQ2). Then we examine the participants’ perception of the explanations and their satisfaction ratings (RQ3). Lastly, both the prediction tasks and the satisfaction survey are followed by reasoning questions that we qualitatively analyze to explore the use of explanation elements in reasoning for the different groups (RQ4).
The first experiment focused on understanding the overall utility of the explanations in understanding and predicting the agent’s actions. This experiment is set in a MiniGrid environment, where we examine the performance of the aggregated version of the explanations in a sparse-reward setting with two reward classes.
In the second study, we evaluate the usefulness and actionability of explanations in a more complex environment. We set our experiment in the Crafter environment to explore the effectiveness of the local explanations generated by our Experiential Explanations method in a more complex environment with sequential multistep goals, with more granular reward classes. Moreover, we want to see if the local Experiential Explanations can convey these complex values to the participants coherently.
In the next sections, we provide a comprehensive overview of the two experiments we designed and conducted to evaluate our explanation technique and address our research questions.
@@ 5 Experiment 1: MiniGrid Environment
In this study, we evaluate our explanations in an environment with two simple reward sources and simplified grid-navigation settings, with only one possible counterfactual action at the decision point (see Figure 3). To answer our research questions we first automatically validated the faithfulness of the explanations (RQ1) and then we evaluated our explanation method with quantitative and qualitative analyses through a human evaluation study (RQ2, RQ3 and RQ4). We built our experimental domain on MiniGrid [29], a popular minimalistic grid world environment that offers a variety of tasks to test RL algorithms.
@@ 5.1 Model Setup and Environment Details
    Negative rewards heavily influence the counterfactual trajectory.
    Template: “If the agent $[a_{user}]$, it will pass through regions influenced by the [negative influence]; going $[a_{agent}]$ feels safer”.
    Explanation Example:“If the agent goes down, it will pass through regions influenced by the dangerous blue obstacles; going up feels safer”.
@@
    Goal rewards have a lower influence on the counterfactual trajectory.
    Template: “If the agent $[a_{user}]$, it will pass through regions less influenced by the [positive influence]; going $[a_{agent}]$ feels better”.
    Explanation Example: “If the agent goes right, it will pass through regions less influenced by the green goal; going left is better”.
@@ 5.2 Baselines
In this experiment, we focus on similar baseline explanations that the agent can generate. We compared our method with the following baselines:
@@
5.3 Faithfulness Evaluation Methodology
To evaluate the reliability of our explanation method, we resort to addressing the faithfulness (RQ1), we combine the predictions of the influence models using two methods. The first is to combine the values of the influence predictors and using their cumulative value to estimate the agent’s action $action=\operatorname*{argmax}_{a}(\sum_{i}P_{i}(s,a)-\sum_{j}N_{j}(s,a))$, where $P_{i}$ is a positive influence predictor and $N_{j}$ is a negative influence predictor. The second method is to train a classifier on the predicted positive and negative influence values as features and the agent’s actions as the ground truth. We used a support vector machine (SVM) classifier with a radial basis function (RBF) kernel. We also hypothesized that in the sparse reward setting, it is more meaningful to look at the influences around the emitting rewards rather than further away from them, as it would likely be more influential to the agent’s decision in those areas rather than further away and therefore, the values would likely be better at predicting the agent’s actions when those influence signals are high enough to impact the agent. To do this, we filter our dataset, by influence values at a set of different fixed intervals and examine the prediction accuracy at each threshold, we start at $0.100$ and we keep increasing by $+0.1$ until it is no longer meaningful to do so. The faithfulness methodology does not make use of the baselines since it is the comparison of the aggregation of information from influence predictors to the main policy model.
5.4 User Study Methodology
@@
We designed our human-participant evaluation following the framework proposed by Hoffman et al. [32], evaluating our system for understandability, performance, and user satisfaction. We conducted our online evaluation between subjects using Prolific [^3] http://prolific.co with 81 participants aged 21-71 years old (M=37.41, SD=11.03); 39.5% of the participants identified as women. The average study completion time was approximately 10 minutes, and we compensated each participant with \$3.75. The study consists of two parts: answering a series of prediction tasks (RQ2, RQ4) and a satisfaction survey (RQ3, RQ4). The explanation group selection were fully randomized and balanced using Qualtrics [^4] https://www.qualtrics.com/.
The study involved participants observing an agent (represented as a red triangle) navigate towards a goal (green square) across three episodes, with certain shapes yielding negative rewards. Initially, participants made uninformed predictions about the agent’s movements, which were manipulated to ensure incorrectness. This established the conditions for which the participant must now learn to correct their understanding of the agent’s interactions with the environment. This setup was used to teach the participants to adjust their understanding based on consistent negative reward assignments in subsequent episodes. The last two episodes were considered for the prediction task. The expectation was that the participants in the experimental condition would demonstrate improved prediction accuracy over those in the baseline conditions. In addition, a satisfaction survey assessed participants’ perceptions of the explanations provided for the agent’s behavior. Each of the questions in the prediction task we followed by a text entry question to capture the participants rationales for making these predictions. In addition, another text entry question was added to get an overall impression of the helpfulness of these explanations. These are the responses that we analyze in our qualitative section. We include screenshots and a detailed description of the study in the Appendix D.
@@ 5.5 Quantitative Results
RQ1: Faithfulness. For measuring faithfulness we look at the accuracy of influence models in predicting agent actions, we found that the accuracy of the prediction of the direct aggregation is $72\%$ and the accuracy of the classifier $76\%$. To examine the alignment of the influence predictors with the agent at states with higher influence, we gradually examine the alignment between the highest action as predicted by the influence predictors and the agent’s actions. At a threshold of $0.100$, the aggregate of influence predictors agreed with the main policy $83\%$, covering at least $20\%$ of the environment. At higher thresholds, agreement increases above $90\%$, indicating that when decisions are the most important, faithfulness is very high. Looking at the positive influence predictor alone, at a threshold of $0.200$ and higher, it alone can predict the agent with a precision of $100\%$. This is to be expected; influence predictors cannot dictate what states and actions get explored—it is controlled by the main policy and the main training algorithm—so states that are less critical to positive reward attainment are under-sampled and prone to more error in both the main policy and the influence predictors. In contrast, more important decision points are sampled more often, so the influence predictors get more training on those states and actions.
RQ2: Prediction Correctness. We measure whether the participant can correctly predict the agent trajectory after interacting with the second and third episodes (Figure 3).
@@
RQ3: Satisfaction Survey. For each type of explanation, we analyzed the ratings of the participants for the Likert scale questions on the satisfaction survey. We include a visualization of the results in Figure 22 in the Appendix. Experiential Explanations had the highest average scores on all dimensions, except trust. A one-way analysis of variance (ANOVA) followed by Tukey’s honestly significant difference test revealed that the Experiential Explanations $(M=3.62,SD=1.32)$ were preferred over the Heatmaps $(M=2.42,SD=1.35)$ in completeness ($F(3,81)=3.032,p<0.05$). Experiential Explanations $(M=3.71,SD=1.0)$ were preferred over the Heatmaps $(M=2.63,SD=1.38)$ in satisfaction($F(3,81)=3.106,p<0.05$). Experiential Explanations $(M=4.38,SD=0.59)$ were preferred over Q-Value Explanations $(M=3.28,SD=1.01)$ and Heatmap Explanations$(M=3.0,SD=1.37)$ in the dimension of understandability ($F(3,81)=5.572,p<0.05$).
5.6 Qualitative Analysis and Findings
@@
To address RQ4, we thematically analyzed open response questions for the tasks and satisfaction survey to better understand how participants used the explanations. Thematic analysis is one of the most common qualitative research methods and involves identifying, analyzing, and reporting patterned responses or meanings (i.e. themes[35]. It is most appropriate for understanding a set of experiences, thoughts, or behaviors across data [36]. To conduct the thematic analysis for each of these codes, three authors performed the following established process, as outlined by [35]: (1) We familiarized ourselves with the data through repeated active readings. (2) We developed a coding framework consisting of a set of codes organizing the data at a granular level and the corresponding well-defined and distinct definitions. (3) Each of the authors independently coded the entirety of the data. If a code was unclear or a new code was needed, the authors consulted with each other to modify the coding framework. If a change was made, each author independently re-analyzed the entirety of the data. (4) We jointly examined the codes to derive themes and findings. We compared our application of the codes for each response to ensure consistency. Discrepancies were counted and used to calculate the inter-rater reliability. The process iterated until all discrepancies were resolved. We identified three themes: Usefulness, Format, and Utilization. The codes, definitions, and examples are summarized in Figures 4 and 5.
Figure 4: Summary tables of the codes used to understand how users across the explanation groups precieved the explanations in terms of usefulness and format.
Figure 5: Summary table of the codes used to understand how users across the explanation groups utilized information from the explanations.
@@
Explanation Usefulness. The theme of Usefulness emerged in the open-ended comments about how participants perceived the explanations. Because these comments were made alongside users’ explanations of their reasoning, this provides additional information on how useful users found the explanations. We identified three Usefulness codes: Useful, Not Useful, or Unclear We coded a response Useful if a participant explicitly mentioned that the explanation was useful, for example, “I understood that purple was dangerous and that the agent should avoid getting near it (EE8).” Responses were coded Not Useful if a participant explicitly stated that the explanation was not useful, for example, “I tried to read the chart but found it confusing and unclear. I didn’t understand it very well (HE10).”. Responses were coded Unclear if the user’s perception was ambiguous, for instance, “I found the explanations a bit too brief. For instance, what makes the purple circles feel safer than the blue squares (EE7)?” The initial inter-rater reliability for Usefulness was 81.5%.
@@
Explanation Format. Participant responses were also coded for Format. We derived three codes for Format: Need More Information, Hard to Understand, and Suggestion. Responses coded as Need More Information explicitly stated why the explanation was lacking and what data was desired. If the response said that the explanation was not clear, it was coded as Hard to Understand, and if the response offered another way to explain, it was coded as Suggestion. Only responses from individuals who received explanations and referenced the explanation’s content could be coded for Format,41.98% of our dataset; therefore, our reported observations are mainly comparisons of the percentages and could not be statistically tested due to the small data size. Participants often expressed a Hard to Understand $40\%$ when shown Heatmap explanations, followed by $29.41\%$ who got Q-value explanations while this was not a problem for the Experiential Explanation group. On the other hand, we observed Need More Information between the three explanation groups with higher concentrations of $35.29\%$ being from the Q-value explanation group and $26.67\%$ are from the Heatmap explanation groups and $21.05\%$ are from the Experiential Explanation group. Participants who did not find the explanations useful had more Format comments, especially requesting more Need More Information $30.77\%$ and Hard to Understand $69.23\%$. The initial inter-rater reliability was 76.47%.
@@
We compared the frequency with which utilization codes appeared in open-responses. If a participant used the same types of information as the explanation provided, we considered the participant as “Utilizing” the explanation. Experiential Explanations provide Object and Reward Related information. Heatmap and Q-Value Explanations provide Reward Related Explanations. No Explanation only provides for Prior Behavior. 71% of the participants who were shown Experiential Explanations utilized the exact kinds of information presented in the explanation, compared to 21% and 28% for the Heatmap and Q-Value explanations, respectively. Although the Q-Value and Heatmap explanations both presented Reward Related information, participants still relied heavily on Prior Behavior to understand the agent.
@@
Figure 6 shows the distribution of the reasons the participants referenced for each type of explanation. Participants who saw Experiential Explanations most used Object related and Reward Related information. Participants who saw Q-Value explanations referenced Goal Related reasons the most. Participants who saw no explanation most often utilized Prior Behavior, followed by Object Related reasons. Notably, participants who saw Experiential Explanations were 4 times less likely to rely on Prior Behavior, indicating they did not rely as much on watching the agent to understand it. Those who saw Heatmap and Q-Value Explanations relied heavily on Prior Behavior and Utilization patterns mimicking those of the No Explanation condition.
Experiential Explanations were 13 and 15 times more likely to have the information they provided be Utilized and Useful, respectively. Utilized refers to what they used in their reasoning, while Useful focuses on whether they thought the explanations were useful or not in the post-task satisfaction survey. Yes, it is the code from the previous section. with respect to Heatmap and Q-Value explanations, respectively. We also found that among those who used components of explanations in their reasoning, 55.88% answered correctly on both questions ($p<0.05,df=1,\chi^{2}=3.849$). But within the Experiential Explanations group, the percentage was 84.62% and that group was more statistically significant to get both questions correctly than other baseline explanation groups ($p<0.05,df=3,\chi^{2}=12.22$).